What are classical limits? Explain how quantum distribution laws are reduced to classical Maxwell-Boltzmann distribution?
1
Expert's answer
2020-03-30T07:58:23-0400

Classical Limits:-The classical limit is the ability of a physical theory to approximate or "recover" classical mechanics when considered over special values of its parameters. The classical limit is used with physical theories that predict non-classical behavior.

The function f(e) depends on whether or not the particles obey the Pauli exclusion principle.

We know that for the bosons distribution function "f_{BE}(\epsilon)=\dfrac{1}{e^{\alpha+\frac{\epsilon}{kT}}-1}"

For fermions, the distribution function is "f_{FE}(\epsilon)=\dfrac{1}{e^{\alpha+\frac{\epsilon}{kT}}+1}"


We know that max well distribution function "f(\epsilon)=Ae^{\frac{-\epsilon}{kT}}"


The number of particles having energy "\epsilon" at temperature T,

"n(\epsilon)= Ag(\epsilon)e^{-\epsilon/kT}"

So, for atom in ground state,

"n(\epsilon_1)= Ag(\epsilon_1)e^{-\epsilon_1/kT}" --------(i)

So, for atom in first excited state,

"n(\epsilon_2)= Ag(\epsilon_2)e^{-\epsilon_21/kT}" -----(ii)

Now dividing equation (ii) by (i)

"\dfrac{n(\epsilon_2)}{n(\epsilon_1)}=\dfrac{g(\epsilon_2)}{g(\epsilon_1)}e^{-\frac{\epsilon_2-\epsilon_1}{kT}}"

="Ae^{-\epsilon/kT}"